A recipe requires 2/3 cup of flour and makes enough for 6 people. how much flour should be used to make enough for 15 people?

Answer by jorel1380(3719)

6/15:2/3/n
6n=15x2/3
6n=10
n=10/6=5/3 cups of flour needed for the recipe..